# Beyond Statistical Estimation: Differentially Private Individual Computation in the Shuffle Model

### v1

@@ VII-B Location-based Social Systems
In this part, we evaluate the performance of PIC model for privacy preserving nearest neighbor (NN) queries on users, a prevalent task in location-based social networks [22]. Like the exemplar spatial crowdsourcing tasks, the nearest location senders’ information (including their corresponding public keys) w.r.t. $\ell_{2}$-distance) of a user is securely return to the user. These matched neighboring users can then securely communicate with other to accomplish subsequent tasks.
TABLE V: Statistics of location social network datasets.
Dataset users check-ins location domain
Gowalla 138368 6442892 $[37.54,37.79]\times[-122.51,-122.38]$
Foursquare 1083 227428 $[40.55,40.99]\times[-74.27,-73.68]$
Datasets and settings. We use use two real check-in datasets, Gowalla dataset [22] and Foursquare dataset [87]. The statistics of two datasets are summarized in Table V. The location domains are rescaled to the standard domain $[-1,1]\times[-1,1]$. As for nearest neighbor queries, we set the nearest radius $\tau$ for Gowalla to $0.1$ and the nearest radius $\tau$ for Foursquare to $0.2$, so that users have a few tens or hundreds neighbors (vary due to check-ins densities). Accordingly, the population size for privacy amplification in PIC model is set to $n-1000$. The concrete cryptography implementations and data randomizers for nearest neighbor queries are the same as the one for spatial crowdsourcing, while the server simply runs a radius-based nearest neighbor algorithm and return the neighboring (public key, noisy location) lists to each user.
Performance metrics. We evaluate the performance of nearest neighboring queries in terms of precision and recall, and $F_{1}$ score of neighboring lists. Let $N_{i}$ denote the true neighbor lists that have $\ell_{2}$ distance to user $i$ less than the parameter $\tau$, and let $\widehat{N}_{i}$ denote the retrieved neighbor lists that have $\ell_{2}$ distance to user $i$ based on noisy (sanitized) locations less than the parameter $\tau$. The precision of nearest neighboring queries is defined as $\text{precision}={(\sum_{i\in[n]}|N_{i}\bigcap\widehat{N}_{i}|)}/{(\sum_{i\in[n]}|N_{i}|)}$, the recall of nearest neighboring queries is $\text{recall}={(\sum_{i\in[n]}|N_{i}\bigcap\widehat{N}_{i}|)}/{(\sum_{i\in[n]}|\widehat{N}_{i}|)}$. We use a balanced indicate
$F_{1}\text{ score}=2\cdot\text{precision}\cdot\text{recall}/(\text{precision}+\text{recall})$
as the major measurement. All presented results are the average value of $10$ experimental runs.
Experimental results. Figure 8 displays the $F_{1}$ scores for radius nearest neighbor queries. To evaluate the effects of varying user numbers, we adjusted the number of users $n$ in the Gowalla dataset from $10000$ to $40000$. Notably, as $n$ surpasses $50000$, the PIC model’s scores see an approximate relative increase of $100\%$ compared to the local model. The Minkowski response consistently records the highest scores. When used in conjunction, they can achieve up to $90\%$ of the $F_{1}$ scores within a reasonable privacy loss (e.g., $\epsilon\leq 3$), especially when $n$ is relatively large.
Figure 8: F1 scores of radius nearest neighbor queries in location-based social systems.
@@ VII-C Federated Learning with Incentives
Recall that in federated learning with rigorous privacy protection, many clients collaboratively contribute gradient updates of model parameters so as to build a machine learning model. Each client $i$ holds a local dataset $D_{i}$. In each iteration $t$, a client $i$ downloads the latest model parameter $W_{t-1}$, and computes the local gradient $g_{i}$ based on local dataset $D_{i}$. Then, the client sanitizes $g_{i}$ and sends it to the shuffler/server. Without loss of generality, we assume the server uses Shapley value [69] to determine the gradient payoffs, where the utility function for computing Shapley value is based on the model parameters’ performance (e.g., accuracy or the negative of loss function) on a validation dataset $D_{val}$. The Shapley values themselves may severely leak sensitive information about user data [59]. Despite the vast non-linear computations involved in evaluating Shapley values with neural networks, no rigorous protection has been offered for federated learning with incentives in existing literature. In this study, we demonstrate the feasibility of computing Shapley value without incurring additional privacy loss.
TABLE VI: Model architecture with 4292 trainable parameters.
@@
Datasets and settings. Our experiments with federated learning use the MNIST handwritten digit dataset, comprising $60,000$ training, $5,000$ validation, and $5,000$ test images. We employ a neural network outlined in Table VI with $d=4292$ trainable parameters. It achieves $98.6\%$ accuracy using non-private, untruncated vanilla SGD and $92.7\%$ accuracy with non-private, truncated SGD. For our experiments, we simulate $50,000$ clients with each client possessing one sample (i.e., $n=50,000$). During each private SGD step, $s=10,000$ clients are randomly selected. Each client subsamples $0.5\%$ of the gradient vector dimensions, then clips the $\ell_{1}$-norm of the block gradient with a threshold of $c=0.001$. We compare our PIC model against a local model, both implementing Laplace mechanism for gradient randomization.
Implementation details. We use Elliptical curve encryption to perform secure point-to-point information delivery. Specifically, to delivering monetary incentives with anonymity protection, clients and server may employ untraceble crypto-currency (e.g., ZCash [46]). Each client creates one random temporary wallet address and append it to the message, the server computes incentive payments of the message and then send corresponding amount of coins to the wallet address.
Performance metrics. We assess performances using gradient/model accuracy and Shapley value accuracy. Gradient accuracy is gauged using the mean squared error (MSE) between the subsampled, clipped, and privatized gradients $\hat{g}_{i}$ and the corresponding non-private gradients ${g}_{i}$:
$\text{gradient}\ MSE=\|\sum\nolimits_{i\in[s]}\hat{g}_{i}-\sum\nolimits_{i\in[s]}{g}_{i}\|_{2}^{2}.$
Model accuracy refers to the image classification accuracy on the test dataset. Shapley value accuracy is the MSE between Shapley values of privatized gradients $\hat{g}_{i}$ and non-private gradients ${g}_{i}$:
$\text{shapley}\ MSE=\sum\nolimits_{i\in[s]}(\widehat{\texttt{Shapley}}_{i}-{\texttt{Shapley}}_{i})^{2},$
where the Shapley values for privatized gradients use the utility function (see Appendix -H for detail):
$\widehat{U}(S)=\frac{<\texttt{grad}_{val},\sum_{i\in S}\hat{g}_{i}>}{\|\texttt{grad}_{val}\|_{2}\cdot\|\sum_{i\in S}\hat{g}_{i}\|_{2}}.$
The presented gradient/shapley MSE averages over $50$ aggregation rounds, and the model accuracy is from the final round.
Experimental results. Since the curator model of DP and other cyrptographic MPC approaches (e.g., secure aggregation [18]) are not applicable to privacy preserving joint gradient aggregation and Shapley value computation (with complicate non-linear computation), we compare equivalent shuffle model with the local model of DP. Both models use Laplace noises [28] to privatizing local gradients, and use tight numerical composition for sequential gradient averaging [42]. We present the experimental results in Table VII, with varying overall privacy budget $\epsilon$. Every result is the average value of $5$ simulations.
TABLE VII: Utility performance of joint gradient aggregation and Shapley-value incentive for federated learning with differential privacy guarantee on MNIST dataset.
@@
MSE
local model $\epsilon=1$ $19.1\%$ $1.7$e-4 $8.9$e-3
$\epsilon=3$ $28.9\%$ $1.3$e-4 $6.1$e-3
$\epsilon=5$ $35.6\%$ $1.1$e-4 $5.2$e-3
PIC model $\epsilon=1$ $77.6\%$ $2.5$e-5 $4.6$e-3
$\epsilon=3$ $81.2\%$ $1.4$e-5 $4.2$e-3
$\epsilon=5$ $85.1\%$ $1.1$e-5 $3.9$e-3
It is observed that PIC model enjoys significantly better gradient accuracy, increase model test accuracy by about $50\%$ in the absolute sense, reduces about $90\%$ gradient aggregation error, when compared to the local model. The PIC model also reduces about $40\%$ error on Shapley values, permit accurate payoff distribution to incentive honest/informative participants.
@@ VIII Conclusion
In this paper, we proposed private individual computation model, for multi-party individual computation with DP preservation, having broad application to spatial crowdsourcing, resource allocation, location-based systems, and federated learning with incentives. The paradigm introduces statistically random identities into the traditional shuffle model of DP, so as to support secure functionalities required by individual computations, meanwhile maintaining the most benefits of privacy amplification via shuffling. It incurs only $O(d)$ computation and communication costs (matching the non-private setting) on the client side, meanwhile permitting plaintext computation on the server side. Additionally, with the increasing of number of participating parties, the utility of individual computation grows (as there are stronger privacy amplification effects), making it suitable for large-scale applications.
The simple paradigm can efficiently solves prevalent tasks that are computationally intractable in MPC approaches (e.g., Ishai’s anonymous model, secure aggregation, or garbled circuits), such as federated deep learning with incentives and large-scale spatial crowdsourcing, necessitating essentially the same overheads as in non-private settings. Moreover, it reduces $40\%$-$90\%$ errors/costs, increases $100\%$-$300\%$ utility, when compared to existing tractable approaches (e.g., local DP).
Discussions and future research. The PIC model guarantees privacy gracefully in cases of malicious clients or collusion, and can adapt to fully decentralized settings without an orchestrating server. Given its alignment with random digital identities in Blockchains and in Ad tracking systems [54], it offers potential to enhance formal privacy in these systems.

### v2

@@ VII-B Location-based Social Systems
TABLE III: Details of location social network datasets.
Dataset check-ins location domain
Gowalla (San Francisco) 138368 $[37.54,37.79]$$\times$$[-122.51,-122.38]$
Foursquare (New York) 227428 $[40.55,40.99]$$\times$$[-74.27,-73.68]$
Datasets We use two real-world datasets: Gowalla dataset [21] and Foursquare dataset [89]. Gowalla and Foursquare are location-based social network websites where users share their locations by checking-in. Details about the two datasets are summarized in Table III. As before, the location data is normalized to $[-1,1]\times[-1,1]$.
LDP randomizers The LDP randomizers used are the same as those in Section VII-A.
Server-side algorithm The server performs the radius-based nearest neighbor (NN) search for the users, which is a common task in location-based social networks [21]. We set the search radius $\tau$ to 0.2, so that each user has several hundreds or thousands of neighbors, varying due to check-in densities. Note that in this application scenario, the returned neighbors may be deanonymized in the post-computation phase. Hence, the actual privacy guarantee in the PIC model depends on the number of users who remain anonymous (see discussion in Section V-C). Considering this, and each user normally has no more than $n\cdot\frac{\pi\tau^{2}}{2^{2}}<n\cdot 3.2\%$ neighbors, we use privacy amplification population $\lfloor n\cdot 90\%\rfloor$ instead of the group size $n$ when calculating the local budget given the global $\epsilon_{c}$.
Figure 8: Expected $\ell_{2}$ distances of reported locations in location-based social systems with the local model of DP.
Figure 9: Expected $\ell_{2}$ distances of reported locations in location-based social systems with PIC model.
Experimental results We first present the single report errors in the LDP model (Fig. 8) and the PIC model (Fig. 9). In our experiments, each check-in report is treated as if it were submitted by a separate user, with all users in the same group. We tested two scenarios: one with a random subset of 10,000 check-ins and the other using all check-ins. On both datasets, PIC demonstrates better utility than LDP, and the Minkowski randomizer consistently performs the best across all settings. Additionally, it is evident that the number of anonymous users is a crucial parameter, significantly impacting utility.
Figure 10: F1 scores of nearest neighbor queries (LDP model).
Figure 11: F1 scores of nearest neighbor queries (PIC model).
The task-specific utility metric we use for nearest neighbor queries is the $F_{1}$ score. Let $N_{i}$ denote the set of true neighbors of user $i$ within an $\ell_{2}$-distance $\tau$, and let $\widehat{N}_{i}$ denote the retrieved neighbor set computed using the sanitized reports. The precision of nearest neighbor queries is defined as $\frac{\sum_{i\in[n]}|N_{i}\cap\widehat{N}_{i}|}{\sum_{i\in[n]}|\widehat{N}_{i}|}$, and the recall is defined as $\frac{\sum_{i\in[n]}|N_{i}\cap\widehat{N}_{i}|}{\sum_{i\in[n]}|N_{i}|}$. The $F_{1}$ score is then calculated as:
$F_{1}\text{ score}=2\cdot\frac{\text{precision}\cdot\text{recall}}{\text{precision}+\text{recall}}.$
In the LDP model (Fig. 10), the $F_{1}$ score is low, even with a large $\epsilon$. While the group size affects the number of neighbors each user has, it does not impact the $F_{1}$ score. Conversely, in the PIC model (Fig. 11), the $F_{1}$ score is significantly higher. With a large group, the Minkowski randomizer can achieve an $F_{1}$ score of 0.9 with a stringent global budget $\epsilon_{c}=1$ and nearly 1 if the budget is loosened to $\epsilon_{c}=3$.
@@ VII-C Federated Learning with Incentives
In this application, the users collaboratively train a model by Federated learning, and the server decides each user’s incentive by how much their local gradient contributes to the global model. To deliver monetary incentive rewards, we assume the use of an untraceable cryptocurrency (e.g., ZCash [47]). Each user can generate an additional public/privacy key pair according to the specification of the cryptocurrency, derive a wallet address from the public key, and append the wallet address to the end of their report. Based on the Shapley values, the server determines the monetary incentives and distributes them to the users using appended wallet addresses.
Datasets We utilize the MNIST handwritten digit dataset to train a simple neural network using federated learning. The MNIST dataset comprises 60,000 images, with 50,000 designated as training samples. Each user receives one training sample and trains a neural network model, as detailed in Table IV, which contains $d=4292$ trainable parameters. In each round, $s=10,000$ users are randomly selected. These selected users locally compute the gradient vector using Stochastic Gradient Descent (SGD), then subsample 0.15% of the gradient vector dimensions (with unsampled dimensions set to zero), and clip the gradient values at a threshold of $c=0.00015$. Each user submits a sanitized version of the subsampled, clipped gradient vector as their report.
TABLE IV: Model architecture of the neural network.
@@
LDP randomizers We use all LDP randomizers as before except for PlanarLaplace, which is unsuitable for multi-dimensional gradient vectors. The randomizers sanitize the non-zero values in the subsampled gradient vectors while leaving the zero values unaffected.
Server-side algorithm The server performs the following steps: first, it takes a sanitized gradient vector and uses randomizer-specific algorithms to estimate the true values of the sampled dimensions. These vectors with estimated values are denoted as $\hat{g}_{i}$. The server then aggregates the $\hat{g}_{i}$ vectors into a global gradient vector using a simple sum-and-average method. This global gradient vector is published so that users can update their local models. Additionally, the server computes the Shapley value for each user, which measures the marginal contribution of each $\hat{g}_{i}$ (see Appendix -H for details).
Experimental results We train each model for 80 rounds using Federated Learning. The utility comparison results are summarized in Tables V and VI, each reflecting a different global privacy budget. In these tables, we present the single report utility (Gradient $\ell_{2}$ error) and the task-specific utility (Shapley $\ell_{2}$ error). The Gradient $\ell_{2}$ error is calculated using the estimated gradients $\hat{g}_{i}$ and the true gradients $g_{i}$ (clipped and with all unsampled dimensions set to 0): $\frac{\sum_{i\in S}\|\hat{g}_{i}-{g}_{i}\|_{2}}{|S|}$, where $S$ is the set of randomly selected users. The Shapley $\ell_{2}$ error is computed as follows:$\frac{\sum\nolimits_{i\in S}\|\widehat{\texttt{Shapley}}_{i}-{\texttt{Shapley}}_{i}\|_{2}}{|S|}$, where $\widehat{\texttt{Shapley}}_{i}$ is derived from $\hat{g}_{i}$ and ${\texttt{Shapley}}_{i}$ is obtained from the unsanitized $g_{i}$. For reference, we also include the final models’ accuracy in the tables.
TABLE V: Utility comparison of federated learning with incentives: global privacy budget $(1,0.01/50000)$.
@@
$\ell_{2}$ Error
local model Staircase [38] $18.29\%$ $15.30$ $0.0586$
Squarewave [58] $13.91\%$ $13.23$ $0.0590$
PrivHS [29] $28.17\%$ $2.44$ $0.0587$
Laplace [31] $21.22\%$ $2.53$ $0.0575$
PrinUnit [12] $23.28\%$ $2.37$ $0.0592$
Minkowski $18.10\%$ $14.45$ $0.0571$
PIC model Staircase [38] $32.51\%$ $0.781$ $0.0563$
Squarewave [58] $33.75\%$ $0.604$ $0.0585$
PrivHS [29] $65.47\%$ $0.267$ $0.0428$
Laplace [31] $72.6\%$ $0.154$ $0.0508$
PrinUnit [12] $74.02\%$ $0.157$ $0.0438$
Minkowski $\mathbf{78.68\%}$ $\mathbf{0.093}$ $\mathbf{0.0363}$
TABLE VI: Utility comparison of federated learning with incentives: global privacy budget $(3,0.01/50000)$.
Setting
Randomization
Mechanism
Test
Accuracy
Gradient
$\ell_{2}$ Error
Shapley
$\ell_{2}$ Error
local model Staircase [38] $19.68\%$ $5.09$ $0.0519$
Squarewave [58] $18.24\%$ $4.34$ $0.0574$
PrivHS [29] $41.52\%$ $0.850$ $0.0561$
Laplace [31] $28.64\%$ $0.845$ $0.0575$
PrinUnit [12] $39.02\%$ $0.803$ $0.0537$
Minkowski $20.73\%$ $3.731$ $0.0545$
PIC model Staircase [38] $44.58\%$ $0.545$ $0.0563$
Squarewave [58] $53.02\%$ $0.407$ $0.0552$
PrivHS [29] $74.49\%$ $0.190$ $0.0383$
Laplace [31] $74.86\%$ $0.104$ $0.0335$
PrinUnit [12] $77.42\%$ $0.098$ $0.0253$
Minkowski $\mathbf{83.43\%}$ $\mathbf{0.055}$ $\mathbf{0.0219}$
We observe in the tables that the utility in the LDP model is significantly worse than in the PIC model, as expected. This is consistent across all metrics and the final model accuracy. In the local model, the performance of Staircase, Squarewave, and Minkowski mechanisms is poorer compared to the others, but for different reasons. Staircase and Squarewave are designed for single-dimensional data, requiring the privacy budget to be split across dimensions when sanitizing vectors, which leads to reduced utility. For Minkowski, the issue lies in the 80-round federated learning process, where the privacy budget for each round is relatively small (e.g., between 0.2 and 0.75). As previously mentioned, the Minkowski mechanism is intended to achieve better utility with a large local privacy budget. In small privacy budget scenarios, it offers no advantage and may even perform worse. In the PIC model, the performance of Staircase and Squarewave remains poor. However, Minkowski now performs the best among all randomizers. This improvement is due to the privacy amplification effect in PIC, which increases the local privacy budget for each round to between 4.0 and 6.0.
@@ VIII Conclusion
Privacy-preserving computation with Differential Privacy (DP) holds great potential for leveraging personal information. While the shuffle model offers a rigorous DP guarantee with enhanced utility, its application is confined to statistical tasks. In this paper, we introduce a novel paradigm called Private Individual Computation (PIC), which extends the shuffle model to scenarios where each user requires personalized outputs from the computation. We demonstrate that PIC can be realized using an efficient protocol that relies on minimal cryptographic operations while maintaining the advantages of privacy amplification through shuffling. To further enhance utility, we developed a local randomizer specifically designed for PIC. We provide formal proofs of the protocol’s security and privacy, as well as the asymptotic optimality of the randomizer. Extensive experiments validate the superiority of the PIC protocol and the randomizer, showcasing their performance across three major application scenarios and various real-world datasets.